An investment of $8000 earned a net gain of $220 in one year. Part of the investment was in bonds and earned income at a rate of 7 %. The rest of the investment was in stocks and lost money at the rate of 10%. How much was invested in bonds?
Answer by checkley77(12569)

-.10x+.07(8000-x)=220
-.10x+560-.07x=220
-.17x=220-560
-.17x=-340
x=-340/-.17
x=$2,000 invested @ 10%
8,000-2,000=6,000 invested @ 7%.
Proof:
-.10*2,000+.07*6,000=220
-200+420=220
220=220